OpenStudy (anonymous):

help me simplify: the answer is supposed to be FX=0.89 FY but i keep getting something else...{(FX (6x12)/(30x10^6) (0.5)} = {FY (4x12)/(15x10^6) (0.75)}

OpenStudy (anonymous):

your copying and pasting doesn't look real pretty

OpenStudy (anonymous):

yeah i know...but that's not the point here...

OpenStudy (anonymous):

well i can't even see what the problem is..... so that's a problem

OpenStudy (anonymous):

{(FX (6x12)/(30x10^6) (0.5)} = {FY (4x12)/(15x10^6) (0.75)}

OpenStudy (anonymous):

i thnk that's much better...

OpenStudy (anonymous):

it's not; need more brackets. unclear what the 0.5 and the 0.75 multiply. you should use the equation button below :)

OpenStudy (whpalmer4):

If the answer given is correct, it should be \[FX\frac{(6*12)}{(30*10^6)(0.5)} = FY\frac{(4*12)}{(15*10^6)(0.75)}\]To simplify, start by striking the common factors (\(12,10^6\)) from both sides: \[FX\frac{(6)}{(30)(0.5)} = FY\frac{(4)}{(15)(0.75)}\]Then multiply both sides by 30: \[FX\frac{(6)}{(30)(0.5)}*30 = FY\frac{(4)}{(15)(0.75)}*30\]\[FX\frac{(6)}{(0.5)} = FY\frac{(4)}{(0.75)}*2\]\[FX (12) = FY\frac{(8)}{(0.75)}\]Divide both sides by 12:\[FX=FY\frac{(8)}{(12)(0.75)} = FY (\frac{8}{9})\]and if you punch 8/9 into your calculator, I think you'll be happy with the result.

OpenStudy (anonymous):

thnks for ur help